Let the sequence Tn be defined by T1 = T2 = T3 = 1 and Tn = Tn-1 + Tn-2 + Tn-3 for n ≥ 4. Use induction to prove that Tn < 2n for n ≥ 4


1
Expert's answer
2021-10-27T13:59:22-0400

Let "P(n)" be the proposition that "T_n<2^n."

Basis Step

"P(4)" is true, because "T_4=T_1+T_2+T_3=1+1+1=3<16=2^4."

Inductive Step

We assume that


"T_{k-1}<2^{k-1}"

"T_{k-2}<2^{k-2}"

"T_{k-3}<2^{k-3}"


"T_k=T_{k-1}+T_{k-2}+T_{k-3}<2^k"

Under this assumption


"T_{k+1}=T_{k}+T_{k-1}+T_{k-2}<2^{k}+2^{k-1}+2^{k-2}"

"=2^{k+1}(\dfrac{1}{2}+\dfrac{1}{4}+\dfrac{1}{8})=2^{k+1}(\dfrac{7}{8})<2^{k+1}"

"P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we

know that "P(n)" is true for all "n\geq4." That is, we have proved that


"T_n=T_{n-1}+T_{n-2}+T_{n-3}<2^n, n\geq 4"